# Is the Top Still Spinning? Evaluating Subjectivity in Narrative Understanding

### v1

@@ 1 Introduction
A possible solution to the problem of factual errors in LLM-generated output lies in having a separate model or process to verify factuality (Durmus et al., 2020; Laban et al., 2022; Chen et al., 2023; Tang et al., 2024). In domains such as mathematical reasoning, verifiers like this can be used not only for evaluation, but at either training time or inference time to improve models (Zelikman et al., 2022; Wang et al., 2024). However, whether an output is factual, or whether it is entailed given some input, has been shown to be highly subjective (Pavlick & Kwiatkowski, 2019; Nie et al., 2020; Jiang & de Marneffe, 2022).This kind of subjectivity is common in tasks in the social sciences and humanities. In this work, we address narrative summarization, which plays a dual role of being a useful application of LLMs in and of itself as well as a proxy task for dealing with complex issues of subjectivity in factual judgments.
@@
To resolve this, we first remove the assumption that there can be a universally agreed upon faithful/unfaithful label for all claims. We instead use LLM-generated rewrites as a method of evaluating claims, which we term the Ambiguity Rewrite Metric (ARM). This method produces a binary judgment (whether or not a claim is rewritten), as well as a correction of the issues in the claim. The degree to which the claim changed also provides signal on how flawed the original claim was. For example, instead of labeling the claim about the spinning top as true or false, a rewrite could specify that the top wobbles and we do not see whether or not it falls. We can quantify the amount of rewriting either through edit distance or the number of eplanation points necessary to justify the changes.
@@ 2 Background
We hypothesize that fixing ambiguities in summary claims can reduce their subjectivity and produce more objectively faithful claims. Approaching evaluation in this way therefore focuses on measuring how far claims are from objective faithfulness rather than trying to assign an objective label up front. We therefore propose using LLM rewrites as an automatic method of evaluation, the Ambiguity Rewrite Metric (ARM). This type of evaluation has been shown useful when executed by humans but we demonstrate that LLM-generated rewrites are an effective evaluation method (Nanba & Okumura, 2004; Liu et al., 2023; yao, ).
@@ 3.1 Claim rewriting
We use a rewrite model to rewrite each claim , producing a rewrite and optionally an explanation of the rewrite consisting of individual points, such that . In practice, we use an LLM as , prompting it to rewrite claims with ambiguities or unfaithful details or just repeat the original claim wording if there are no issues.
@@ 4.4 Baselines and rewrite models
We also compare against a fine-tuned model. We use a Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) and the synthetically generated data of objective and subjective claims discussed in Section 4.1. This method results in a dataset of about 2k claims that we can finetune on. We apply LoRA (Hu et al., 2022) with a rank of 64, alpha set to 64, and a dropout rate of 0.05. For hyperparameters, we use a learning rate of 5e-5, a batch size of 8, gradient accumulation over 2 steps, and train for 1 epoch.
@@ 5.2 Can rewrites detect subjectivity?
| Method | Balanced accuracy | F1-macro |
@@
The ARM metric labels claims which get rewritten as subjective and labels them objective otherwise. We compare how well ARM detects subjectivity relative to the baselines detailed in Section 4.4 in Table 2. We find that finetuning a detection model on the synthetic dataset does not transfer effectively to the original claims. In zero-shot and few-shot settings, we see that Claude is overall a stronger model for detecting subjectivity than GPT-4. Self-consistency does not work well for Claude but has comparable performance to other methods for GPT-4. ARM is a strong method for detecting subjectivity across both models, and the strongest when looking at balanced accuracy. Overall, ARM is a stronger or comparable method to other baselines for binary evaluation while providing the added benefits of the rewrite itself and explanation. We report additional results with alternate prompts and against faithfulness labels in Appendix A.7 but observe that ARM is fairly consistent across different prompts.
@@ 6 Related Work
Ambiguity in evaluation Ambiguity has been studied in natural language entailment (Pavlick & Kwiatkowski, 2019; Nie et al., 2020; Jiang & de Marneffe, 2022), a related task to summary faithfulness evaluation. Other work has studied how to improve or manage annotator disagreement (Uma et al., 2022; Krishna et al., 2023; Min et al., 2023). Most similar to ours is Koupaee et al. (2025), which examines ambiguity in evaluating meeting summaries. Their taxonomy of ambiguities supports our findings and by working with narratives, we incorporate the added challenge of intentional ambiguity in the source text.
@@ 7 Conclusion
We follow protocol approved by our IRB for the human annotation work in this study. One of the authors holds an equity interest in OpenAI.
@@ A.6 Additional ambiguity analysis details
For the inductive thematic analysis, the two authors involved write an explanation for the ambiguities involved in each subjective claim. They then discuss these explanations and arrive at four types of ambiguities. They go back and code each subjective claim with one of these types and discuss and adjudicate any disagreements.
@@
A.7 Does the rewrite model and prompt affect results?
| Claude subj. | 67.89 | 67.42 | 69.00 | 246 | 0.40 |
| Claude both | 69.15 | 68.05 | 69.96 | 341 | 0.61 |
| Claude inconsist. | 66.7 | 68.21 | 63.56 | 218 | 0.46 |
| GPT-4 subj. | 53.95 | 55.42 | 54.61 | 486 | 0.17 |
| GPT-4 both | 63.55 | 64.95 | 63.39 | 203 | 0.22 |
| GPT-4 inconsist. | 64.01 | 63.09 | 63.56 | 390 | 0.21 |
We assess how sensitive rewrites are to prompting method and model choice in Table 5. We evaluate detection against subjectivity labels, faithfulness labels, and claims that are subjective or unfaithful. We use the rewriting prompts shown in Appendix A.2. We once see that Claude is a stronger model for rewriting in terms of binary classification of subjectivity or faithfulness. We see that GPT-4 provides the benefit of making more minimal changes to claims though as its rewrites have less than half the edit distance relative to Claude. Overall, the LLM used for rewriting matters more than the prompt with the exception of the subjectivity-targeted prompt for GPT-4. This setting produces the most rewrites by far which leads to its poor performance.

### v2

@@ Abstract
Is the Top Still Spinning? Evaluating Subjectivity
in Narrative Understanding
Melanie Subbiah1, Akankshya Mishra1, Grace Kim2, Liyan Tang2, Greg Durrett2, Kathleen McKeown1 1Columbia University, 2The University of Texas at Austin Correspondence: m.subbiah@columbia.edu
@@ 1 Introduction
A possible solution to the problem of factual errors in LLM-generated output lies in having a separate model or process to verify factuality Durmus et al. (2020); Laban et al. (2022); Chen et al. (2023); Tang et al. (2024). In domains such as mathematical reasoning, verifiers like this can be used not only for evaluation, but at either training time or inference time to improve models (Zelikman et al., 2022; Wang et al., 2024). However, whether an output is factual, or whether it is entailed given some input, has been shown to be highly subjective (Pavlick and Kwiatkowski, 2019; Nie et al., 2020; Jiang and de Marneffe, 2022).This kind of subjectivity is common in tasks in the social sciences and humanities. In this work, we address narrative summarization, which plays a dual role of being a useful application of LLMs in and of itself as well as a proxy task for dealing with complex issues of subjectivity in factual judgments.
@@
To resolve this, we first remove the assumption that there can be a universally agreed upon faithful/unfaithful label for all claims. We instead use LLM-generated rewrites as a method of evaluating claims, which we term the Ambiguity Rewrite Metric (ARM). This method produces a binary judgment (whether or not a claim is rewritten), as well as a correction of the issues in the claim. The degree to which the claim changed also provides signal on how flawed the original claim was. For example, instead of labeling the claim about the spinning top as true or false, a rewrite could specify that the top wobbles and we do not see whether or not it falls. We can quantify the amount of rewriting either through edit distance or the number of explanation points necessary to justify the changes.
@@ 2 Background
We hypothesize that fixing ambiguities in summary claims can reduce their subjectivity and produce more objectively faithful claims. Approaching evaluation in this way therefore focuses on measuring how far claims are from objective faithfulness rather than trying to assign an objective label up front. We therefore propose using LLM rewrites as an automatic method of evaluation, the Ambiguity Rewrite Metric (ARM). This type of evaluation has been shown useful when executed by humans but we demonstrate that LLM-generated rewrites are an effective evaluation method (Nanba and Okumura, 2004; Liu et al., 2023; Yao et al., 2023).
@@ 3.1 Claim rewriting
We use a rewrite model to rewrite each claim , producing a rewrite and optionally an explanation of the rewrite consisting of individual points, such that . In practice, we use an LLM as (with ), prompting it to rewrite claims with ambiguities or unfaithful details or just repeat the original claim wording if there are no issues.
@@ 4.4 Baselines and rewrite models
We also compare against a fine-tuned model. We use a Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) and the synthetically generated data of objective and subjective claims discussed in Section 4.1. This method results in a dataset of about 2k claims that we can finetune on. We apply LoRA (Hu et al., 2022) with a rank of 64, alpha set to 64, and a dropout rate of 0.05. For hyperparameters, we use a learning rate of 5e-5, a batch size of 8, gradient accumulation over 2 steps, and train for 1 epoch using two A100 GPUs.
@@ 5.2 Can rewrites detect subjectivity?
| Method | Bal. acc. | F1-macro |
@@
The ARM metric labels claims which get rewritten as subjective and labels them objective otherwise. We compare how well ARM detects subjectivity relative to the baselines detailed in Section 4.4 in Table 2. We find that finetuning a detection model on the synthetic dataset does not transfer effectively to the original claims. In zero-shot and few-shot settings, we see that Claude is overall a stronger model for detecting subjectivity than GPT-4. Self-consistency does not work well for Claude but has comparable performance to other methods for GPT-4. ARM is a strong method for detecting subjectivity across both models, and the strongest when looking at balanced accuracy. Overall, ARM is a stronger or comparable method to other baselines for binary evaluation while providing the added benefits of the rewrite itself and explanation. We report additional results with alternate prompts and against faithfulness labels next and observe that ARM is fairly consistent across different prompts.
5.3 Does the rewrite model and prompt affect results?
| Claude subj. | 67.89 | 67.42 | 69.00 | 246 | 0.40 |
| Claude both | 69.15 | 68.05 | 69.96 | 341 | 0.61 |
| Claude unfaith. | 66.7 | 68.21 | 63.56 | 218 | 0.46 |
| GPT-4 subj. | 53.95 | 55.42 | 54.61 | 486 | 0.17 |
| GPT-4 both | 63.55 | 64.95 | 63.39 | 203 | 0.22 |
| GPT-4 unfaith. | 64.01 | 63.09 | 63.56 | 390 | 0.21 |
We assess how sensitive rewrites are to prompting method and model choice in Table 3. We evaluate detection against subjectivity labels, faithfulness labels, and claims that are subjective or unfaithful. We use the rewriting prompts shown in Appendix A.2. We see that Claude is a stronger model for rewriting in terms of binary classification of subjectivity or faithfulness. GPT-4 provides the benefit of making more minimal changes to claims as its rewrites have less than half the edit distance relative to Claude. Overall, the LLM used for rewriting matters more than the prompt with the exception of the subjectivity-targeted prompt for GPT-4. This setting produces the most rewrites by far which leads to its poor performance.
@@ 6 Related Work
Ambiguity in evaluation Ambiguity has been studied in natural language entailment (Pavlick and Kwiatkowski, 2019; Nie et al., 2020; Jiang and de Marneffe, 2022), a related task to summary faithfulness evaluation. Other work has studied how to improve or manage annotator disagreement (Uma et al., 2022; Krishna et al., 2023; Min et al., 2023). Most similar to ours are Koupaee et al. (2025); Mishra et al. (2024); Ramprasad et al. (2024), which include various types of ambiguities in evaluating meeting summaries, real-world QA, and dialogue summaries respectively. Their taxonomies of ambiguities support our findings and by working with narratives, we incorporate the added challenge of intentional ambiguity in the source text.
@@ 7 Conclusion
There are no risks involved with this work as it explores summarization evaluation with publicly available data. We follow protocol approved by Columbia University IRB protocol AAAS4051 for the human annotation work in this study. We use AI assistants to answer questions about coding for this work, but do not use them for any part of research design or paper writing. The StorySumm dataset is available for research use and our use is in line with this purpose. One of the authors, Melanie Subbiah, holds an equity interest in OpenAI.
Limitations
The StorySumm dataset is relatively small which limits the generalization of our conclusions. However, we are still able to see statistically significant results with this dataset. Using human evaluation is a thorough and rich source of feedback but can limit reproducibility of results since different annotators will perform the task slightly differently. Finally, given that we study subjectivity in this work, there is inherent subjectivity involved in the task which can further limit the reproducibility of results. Despite this, we feel it is important to try to formalize and make progress on areas of evaluation that contain grey areas.
Acknowledgments
We would like to express our gratitude to the Upwork workers who contributed annotations for this work. Additionally, we would like to thank our reviewers for their thoughtful feedback. This work is supported by the funds provided by several organizations: the Columbia Amazon CAIT PhD Fellowship, Northrup Grumman, the National Science Foundation, DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence), and Good Systems,444https://goodsystems.utexas.edu/ a UT Austin Grand Challenge to develop responsible AI technologies.
@@ A.6 Additional ambiguity analysis details
For the inductive thematic analysis, the two authors involved write an explanation for the ambiguities involved in each subjective claim. They then discuss these explanations and arrive at four types of ambiguities. They go back and code each subjective claim with one of these types and discuss and adjudicate any disagreements. We observed agreement between the authors on 89% of the sentences (0.53 Cohen’s Kappa) prior to adjudication.